likeabossssssss:

what is this 7.2+(−5.5)=−5.5+7.2?

KjSaif:

7.2−5.5 = −5.5+7.2 1.7=1.7

KjSaif:

Your answer is 1.7=1.7

likeabossssssss:

thx lol

umer:

mhm

MsBrains:

gg